The sum of three consecutive odd integers is 41 more than twice the smallest.Find the three integers

Answer by checkley79(3341)

LET X, X+2 & X+4 BE THE 3 ODD NUMBERS.
X+X+2+X+4=41+2X
3X+6=41+2X
3X-2X=41-6
X=35 ANS. FOR THE FIST INTEGER.
35+2=37 ANS. FOR THE MIDDLE INTEGER.
35+4=39 ANS. FOR THE LAST INTEGER.
PROOF:
35+37+39=41+2*35
111=41+70
111=111